OpenStudy (anonymous):

1/6^x+2=216 How do you solve for x?

OpenStudy (anonymous):

do I understrand the equation right? \[\frac {1}{6^x+2}=216 \]

OpenStudy (anonymous):

log?

OpenStudy (anonymous):

it's and expodentail x+2 instead of 6^x and then add 2 to all of it. And we haven't been taught the log method yet.

OpenStudy (anonymous):

so this? \[\frac {1}{6^{x+2}}=216\]

OpenStudy (anonymous):

\[\frac {1} {6^x} + 2 = 216\]

OpenStudy (ranga):

I think it is: \[\Large \frac{ 1 }{ 6^{x+2} } = 216\]

OpenStudy (anonymous):

\[\left( 1/6 \right)^{x+2}=216\]

OpenStudy (ranga):

216 = 6^3

OpenStudy (anonymous):

3 different ways it can be read...and the 'question asker' isn't even here...

OpenStudy (ranga):

6^(-x-2) = 6^3 -x-2 = 3 x = -5